# A Review of Gravitational Memory and BMS Frame Fixing in Numerical Relativity

### v1

@@ I Introduction
Furthermore, even if one uses perturbation theory to try to predict the GWs, this fails to produce reliable results during the merger phase of the binary, which is typically the loudest and most detectable part of the GW signal. So, even GW models that are built using perturbation theory or by making certain phenomenological assumptions, like EOB or Phenom models Husa et al. (2016); Khan et al. (2016); Pratten et al. (2021); Estellés et al. (2022); Taracchini et al. (2012); Lackey et al. (2014); Pan et al. (2014); Bohé et al. (2017); Ossokine et al. (2020); Ramos-Buades et al. (2023), need to be calibrated against NR waveforms. Consequently, NR simulations, which can achieve arbitrary accuracy with the right computational tools, are at the heart of producing accurate and robust solutions to GR.
@@
Nonetheless, NR can still fail to accurately simulate GR if the code infrastructure is not formulated correctly or if the necessary numerical resolution is not achieved. One such example of this inaccuracy was the inability of NR simulations to resolve a collection of observables in GR that are colloquially referred to as memory effects. These effects are not-yet observed, nonlinear predictions of GR that physically correspond to the net displacement that two freely-falling observers will experience due to the passage of transient GWs. However, apart from being a curious prediction of GR, what makes memory effects particularly tantalizing is that they are intimately tied—through conservation laws—to the symmetry group of future null infinity—part of the asymptotic boundary of spacetime. This symmetry group is not the usual Poincaré group of special relativity, but is a larger group called the BMS group Bondi et al. (1962); Sachs (1962a). Thus, there is hope that with the detection of memory we can not only conduct more stringent tests of GR, but we may even obtain a better understanding of the asymptotic structure of our encompassing universe, which is of immense interest to theorists trying to formulate a theory of quantum gravity through topics such as celestial holography Strominger (2017); Pasterski (2019); Raclariu (2021); Pasterski (2021); Pasterski et al. (2021).
@@
In this review, we highlight recent advancements made in the NR community to resolve memory effects as well as some recent work showing how fixing the BMS freedoms at future null infinity drastically improves both the accuracy and robustness of GW models and analyses. Specifically, in Sec. II we begin by providing some motivation for and intuition behind the BMS group and memory effects. Next, in Sec. III we provide a review of the literature surrounding memory, the BMS group, and BMS frame fixing. Then, in Sec. IV, we provide a more formal explanation of the origins of the BMS group and how memory effects can be understood as stemming from certain conservation laws related to the symmetries of null infinity. In Sec. V, we transition to a review of the code frameworks used to compute gravitational waves at null infinity and we highlight the advancements in the NR community that has made the resolution of memory effects possible. Then, in Sec. VI, we demonstrate the formalism from Sec. IV using binary black hole merger simulations. In particular, we present how the BMS conservation laws can be used to efficiently analyze gravitational waves and understand memory effects. Furthermore, in this section we also provide a review of the detectability of memory and the forecast for its detection in the coming decade. In Sec. VII we highlight the often-overlooked importance of fixing the BMS freedom of NR waveforms to ensure that waveform modeling is performed both accurately and robustly. We show this by reviewing the notion of a superrest frame Mitman et al. (2021a); Magaña Zertuche et al. (2022); Mitman et al. (2022), and demonstrating its utility by comparing NR waveforms to post-Newtonian waveforms and by fitting NR waveforms with predictions made by black hole perturbation theory. Finally, in Sec. VIII we summarize the main points of this review and provide some outlook regarding the future of numerical relativity, memory effects, and testing the nature of gravity.
II Pedagogical approach to BMS and memory
@@ II.1 Motivation
As an alternative perspective regarding the subtleties of coordinates in GR, consider a numerical simulation. At the simplest level, numerical relativists must produce waveforms as tables of timestamps and corresponding strain values measured at a variety of angular locations encompassing the source. [^3] In reality, numerical relativists provide the angular dependence of the strain by representing the strain with respect to some set of angular basis functions, like spin-weighted spherical harmonics. However, to make the connection to coordinates more apparent, in this discussion we ignore this detail and instead consider the naive representation in terms of points on the two-sphere. However, the meaning of those time coordinates, the angular locations at which the strain is measured, and the basis with respect to which the strain is evaluated rely on the essentially arbitrary coordinates used in the numerical simulation—coordinates imprinted by the vagaries of initial data and complicated gauges. Despite confusing declarations that may be found in the NR literature, no GW extraction method can produce “invariant” results in the sense of being independent of the choice of coordinates. Even at linear order in the size of the gauge perturbation, every waveform description is coordinate-dependent. But this issue is not unique to NR simulations. Other gravitational wave modelers, such as those working in post-Newtonian (PN) or even post-Minkowskian (PM) theory, face similar ambiguities. Ideally, we would resolve these coordinate issues in a consistent way, so that waveforms from other simulations, or other models, can be compared to each other.
@@ II.2 Bondi gauge and inertial observers
Now, apart from constructing a coordinate system and a region of spacetime that can be used to study GWs that agrees (up to $1/r$ corrections) with what a distant, inertial observer would see, the other import consequence of constructing the Bondi frame is that, by doing so, we have drastically reduced the coordinate ambiguity that would otherwise plague our GW waveforms. In particular, for the numerical simulation example, we no longer have to express the waveform in terms of the simulation’s arbitrary coordinates, which are ambiguous up to the entire diffeomorphism group. Nor do we have to integrate a family of timelike geodesics and evaluate the waveform along those curves, still ambiguous up to the choice of initial conditions for each geodesic. Instead, we can simply express the waveform as a function of $(u,\theta,\phi)$ on $\mathcal{I}^{+}$.
But one obvious question persists: how much ambiguity still remains in this Bondi frame description? Intuitively, we might expect that the answer would simply be the transformations that leave the usual Minkowski metric unchanged, i.e., the Poincaré group. In fact, this is indeed what Bondi, van der Burg, Metzner, and Sachs thought they would find when studying the symmetry group of future null infinity. However, this intuition turns out to be only nearly correct. Specifically, the symmetry group of future null infinity is the Poincaré group, but with the usual spacetime translations replaced by a larger set of spacetime transformations called “supertranslations”. This is the BMS group.
@@ II.3 The BMS group
with an arbitrary (sufficiently smooth) function $\alpha(\theta,\phi)$, we can check that the asymptotic form of the metric does not change under such a transformation. The function $\alpha(\theta,\phi)$ is exactly a supertranslation and is the ingredient that is needed for constructing the BMS group. It contains the usual spacetime translations as $\ell=0,1$ components, when viewed in terms of spherical harmonic modes.
Like the usual spacetime translations, we may combine two supertranslations by pointwise addition, and can thus turn the set of supertranslations into an abelian group $\mathbb{T}$. In fact, this abelian group $\mathbb{T}$ is a normal subgroup of the full BMS group, with the factor group of the BMS group by $\mathbb{T}$ being the usual restricted Lorentz group $\mathrm{SO}^{+}(3,1)$. The latter, however, is not a normal subgroup. Therefore, the full BMS group is formally the semidirect product $\mathbb{T}\rtimes\mathrm{SO}^{+}(3,1)$. Put more simply, we can express any BMS transformation as a supertranslation followed by a Lorentz transformation (see Appendix A for more details). For a boost with velocity $\vec{v}$, and thus a conformal factor
@@
the combined effect on the retarded time of a supertranslation $\alpha(\theta,\phi)$ followed by this Lorentz transformation is simply Eq. (3) but with a Doppler factor, i.e.,
@@ II.4 Understanding supertranslations
For the moment, consider a pure supertranslation, i.e., a supertranslation which is not a time or space translation. In terms of spherical harmonics, this corresponds to a function $\alpha(\theta,\phi)$ whose spherical harmonics decomposition only consists of $\ell\geq 2$ modes. Formally, supertranslations are rather elementary: they are angle-dependent offsets in the retarded time, as illustrated through Eq. (3). Specifically, for each distant inertial observer at $\mathcal{I}^{+}$, a supertranslation corresponds to a simple change in the origin of the time coordinate. To understand why they are symmetries of future null infinity, we consider the following thought experiment.
Consider some asymptotically flat spacetime with an isolated astrophysical event—like a supernovae explosion or a binary black hole merger—that is emitting radiation, such as photons or gravitational waves, outward in a spherical manner. Furthermore, imagine a network of distant inertial observers that are surrounding this event, each at some finite radius from the event that need not be the same as the other observers. If these observers can communicate, then they could—in principle—use their knowledge of their locations relative to each other and to the central event to synchronize their clocks to ensure that their measurements of the central event are simultaneous in some sense. Now, consider what happens if these observers were placed at larger (but finite) radii. At these new, farther away positions, the signals that they use to synchronize their clocks will take longer to travel between them, but there is no fundamental obstacle to this synchronization in principle.
@@ II.5 The effects of BMS transformations
Figure 1: “Cylinder” diagram of $\mathcal{I}^{+}$ to provide intuition about (i) boosts, (ii) space translations, and (iii) proper supertranslations. The retarded time coordinate $u$ runs vertically in each plot, while the polar coordinate $\theta$ runs azimuthally. Black circles correspond to surfaces of constant $u$ in the untransformed frame. A Lorentz boost dilates $u$ by some factor at each $\theta$ point, while a space translation and a proper supertranslation instead shift $u$ by some function of $\theta$, which can be written as some combination of $\ell=1$ spherical harmonics for space translations, and some combination of $\ell\geq 2$ spherical harmonics for proper supertranslations. The boost is in the $\hat{z}$ direction; the space translation is proportional to the spherical harmoinc $Y_{(1,0)}(\theta,\phi)$, i.e., the $\hat{z}$ direction; and the proper supertranslation is proportional the spherical harmonic $Y_{(2,0)}(\theta,\phi)$.
@@
From this coordinate transformation, by examining how the tetrad on $\mathcal{I}^{+}$ transforms, one can then ascertain that the gravitational wave shear $\sigma$, which is related to the strain $h$ via $h\equiv 2\bar{\sigma}$, and the Weyl scalars transform as
@@
With this information, one then has everything that is needed to transform asymptotic data on $\mathcal{I}^{+}$ and therefore fix the coordinate freedom of that data to match some canonical frame. This notion of mapping asymptotic data to a certain frame is called BMS frame fixing and will be reviewed in Sec. VII.
@@ II.6 Pedagogical approach to memory
This expression, in fact, is nearly correct. The one piece of information that is missing is that these two terms on the right-hand side of Eq. (II.6) need not fully cancel out. One way to realize this is by considering the scattering of two particles in linearized gravity Kovacs and Thorne (1978). First note that, because of the linearization, there will be no energy flux. However, because the particles scatter, there will still be a change in the angle-dependent mass, i.e., a change in the mass multipole moment. Then, because a change in the mass multipole moment corresponds to a change in the shear—or, equivalently, the strain—one can intuit that there should also be a term on the left-hand side of Eq. (II.6) that corresponds to the shear, e.g.,
@@
Using Eq. (II.6), one can then learn a large amount of interesting physics about gravitational waves. For one, we observe that, apart from the gravitational wave strain being sourced by the mass multipole moment, there is also a contribution from an angle-dependent energy flux. Furthermore, if we think about the net change in the strain between non-radiative regimes, i.e., the memory, we immediately find that there are two contributions: one from the angle-dependent mass and one from the angle-dependent energy. These unique contributions are the ordinary and null contributions to the memory Bieri and Garfinkle (2014). [^9] Previously, these contributions were called linear and nonlinear (or Christodoulou). However, they have since been renamed to better reflect the way in which they are sourced Zel’dovich and Polnarev (1974); Braginsky and Grishchuk (1985); Braginsky and Thorne (1987); Christodoulou (1991); Thorne (1992). They can be understood as being sourced by two types of physical processes. In particular, ordinary memory is sourced by systems that contain unbound matter, i.e., massive particles, which approach future timelike infinity. Examples of systems that source this effect are hyperbolic black hole encounters, supernovae, or neutron star mergers. Meanwhile, null memory is sourced by null radiation that escapes to null infinity, e.g., gravitational waves or electromagnetic waves. For an example of what the null memory looks like in the GW produced by a binary black hole merger, see Fig. 2.

### v2

@@ I Introduction
Furthermore, even if one uses perturbation theory to try to predict the GWs, this fails to produce reliable results during the merger phase of the binary, which is typically the loudest and most detectable part of the GW signal. Consequently, GW models that are built by using perturbation theory or GW phenomenology, like effective one-body (EOB) or Phenomenological (Phenom) GW models Buonanno and Damour (1999); Husa et al. (2016); Khan et al. (2016); Pratten et al. (2021); Estellés et al. (2022); Taracchini et al. (2012); Lackey et al. (2014); Pan et al. (2014); Bohé et al. (2017); Ossokine et al. (2020); Ramos-Buades et al. (2023), always need to be calibrated against NR waveforms. Thus, NR simulations, which can achieve arbitrary accuracy with the right computational tools, are at the heart of producing accurate and robust solutions for GW-emitting phenomena in GR.
@@
Nonetheless, NR can still fail to accurately simulate GR if the code infrastructure is not formulated correctly or if the necessary numerical resolution is not achieved. One such example of this inaccuracy was the inability of NR simulations to resolve a collection of observables in GR colloquially referred to as memory effects Zel’dovich and Polnarev (1974); Braginsky and Grishchuk (1985); Favata (2010). These effects are not-yet observed, nonlinear predictions of GR that physically correspond to the net displacement that two freely-falling observers will experience due to the passage of transient GWs. However, apart from being a curious prediction of GR, what makes memory effects particularly tantalizing is that they are intimately tied—through conservation laws—to the symmetry group of future null infinity—part of the asymptotic boundary of spacetime. This symmetry group is not the usual Poincaré group of special relativity, but is a larger group called the BMS group, named after Bondi, van der Burg, Metzner, and Sachs Bondi et al. (1962); Sachs (1962a). Thus, there is hope that with the detection of memory we can not only conduct more stringent tests of GR, but we may even obtain a better understanding of the asymptotic structure of our encompassing universe, which is of immense interest to theorists trying to formulate a theory of quantum gravity through topics such as celestial holography Strominger (2017); Pasterski (2019); Raclariu (2021); Pasterski (2021); Pasterski et al. (2021).
@@
In this review, we highlight recent advancements made in the NR community to resolve memory effects as well as some work showing how fixing the BMS freedoms at future null infinity drastically improves both the accuracy and robustness of GW models and analyses. Specifically, in Sec. II we begin by providing some motivation for and intuition behind the BMS group and memory effects. Next, in Sec. III we provide a review of the literature on memory, the BMS group, and BMS frame fixing. Then, in Sec. IV, we provide a more formal explanation of the origins of the BMS group and how memory effects can be understood as stemming from certain conservation laws related to the symmetries of null infinity. In Sec. V, we transition to a review of the code frameworks used to compute gravitational waves at null infinity and we highlight the advancements in the NR community that has made the resolution of memory effects possible. Then, in Sec. VI, we demonstrate the formalism from Sec. IV using binary black hole merger simulations. In particular, we present how the BMS conservation laws can be used to efficiently analyze gravitational waves and understand memory effects. Furthermore, in this section we also provide a review of the detectability of memory and the forecast for its detection in the coming decade. In Sec. VII we highlight the often-overlooked importance of fixing the BMS freedom of NR waveforms to ensure that modeling of NR waveforms is performed accurately and robustly. We show this by introducing the superrest frame, and demonstrating its utility by comparing NR waveforms to post-Newtonian waveforms and by fitting NR waveforms with predictions made by black hole perturbation theory. Finally, in Sec. VIII we summarize the main points of this review and provide some outlook regarding the future of numerical relativity, memory effects, and testing the nature of gravity with gravitational waves.
II Pedagogical approach to
BMS and memory
@@ II.1 Motivation
As an alternative perspective regarding the subtleties of coordinates in GR, consider a numerical simulation. At the simplest level, numerical relativists must produce waveforms as tables of timestamps and corresponding GW strain values measured at various angular locations encompassing the source. [^3] In reality, numerical relativists provide the angular dependence of the strain by representing the strain with respect to some set of angular basis functions, like spin-weighted spherical harmonics. However, to make the connection to coordinates more apparent, in this discussion we ignore this detail and instead consider the naive representation in terms of points on the two-sphere. However, the meaning of those time coordinates, the angular locations at which the strain is measured, and the basis with respect to which the strain is evaluated rely on the essentially arbitrary coordinates used in the numerical simulation—coordinates imprinted by the vagaries of initial data and complicated gauges. Despite confusing declarations that may be found in the NR literature, no GW extraction method can produce “invariant” results in the sense of being independent of the choice of coordinates Lehner and Moreschi (2007). Even at linear order in the size of the gauge perturbation, every waveform description is coordinate-dependent. But this issue is not unique to NR simulations. Other gravitational wave modelers, such as those working in post-Newtonian (PN) or even post-Minkowskian (PM) theory, face similar ambiguities. Ideally, we would resolve these coordinate issues in a consistent way, so that waveforms from other simulations, or other models, can be compared to each other.
@@ II.2 Bondi gauge and inertial observers
Now, apart from constructing a coordinate system and a region of spacetime that can be used to study GWs that agree (up to $1/r$ corrections) with what a distant, inertial observer would see, the other import consequence of constructing the Bondi frame is that, by doing so, we have drastically reduced the coordinate ambiguity that would otherwise plague our GW waveforms. In particular, for the numerical simulation example, we no longer have to express the waveform in terms of the simulation’s arbitrary coordinates, which are ambiguous up to the entire diffeomorphism group. Nor do we have to integrate a family of timelike geodesics and evaluate the waveform along those curves, still ambiguous up to the choice of initial conditions for each geodesic. Instead, we can simply express the waveform as a function of $(u,\theta,\phi)$ on $\mathcal{I}^{+}$.
But one obvious question persists: how much ambiguity still remains in this Bondi frame description? Intuitively, we might expect that the answer would simply be the transformations that leave the usual Minkowski metric unchanged, i.e., the Poincaré group. In fact, this is indeed what Bondi, van der Burg, Metzner, and Sachs thought they would find when studying the symmetry group of future null infinity. However, this intuition turns out to be only nearly correct. Specifically, the symmetry group of future null infinity is the Poincaré group, but with the usual four spacetime translations replaced by a larger set of spacetime transformations called supertranslations. This is the BMS group.
@@ II.3 The BMS group
with an arbitrary (sufficiently smooth) function $\alpha(\theta,\phi)$, we can check that the asymptotic form of the metric does not change under such a transformation. The function $\alpha(\theta,\phi)$ is exactly a supertranslation and is the ingredient that is needed for constructing the BMS group. It contains the usual spacetime translations as $\ell=0,1$ components, when viewed in terms of spherical harmonic modes, and proper supertranslations as $\ell\geq 2$ components.
Like the usual spacetime translations, we may combine two supertranslations by pointwise addition, and can thus turn the set of supertranslations into an abelian group $\mathbb{T}$. In fact, this abelian group $\mathbb{T}$ is a normal subgroup of the full BMS group, with the factor group of the BMS group by $\mathbb{T}$ being the usual restricted Lorentz group $\mathrm{SO}^{+}(3,1)$. The latter, however, is not a normal subgroup. Therefore, the full BMS group is formally the semidirect product $\mathbb{T}\rtimes\mathrm{SO}^{+}(3,1)$. Put more simply, we can express any BMS transformation as a supertranslation followed by a Lorentz transformation (see Appendix A for more details). For a boost with velocity $\vec{v}$, and thus a conformal factor [^8] See Appendix D of Ref. Boyle (2016) for more on why boosts induce a conformal transformation of the celestial sphere.
@@
where $\hat{n}(\theta,\phi)$ is the unit normal vector to the point $(\theta,\phi)$, the effect on the retarded time of a supertranslation $\alpha(\theta,\phi)$ followed by this Lorentz transformation is simply Eq. (3) but with a Doppler factor, i.e.,
@@ II.4 Understanding supertranslations
For the moment, consider a proper supertranslation, i.e., a supertranslation which is not a time or space translation. In terms of spherical harmonics, this corresponds to a function $\alpha(\theta,\phi)$ whose spherical harmonic decomposition only consists of $\ell\geq 2$ modes. Formally, supertranslations are rather elementary: they are angle-dependent offsets in the retarded time, as illustrated through Eq. (3). Specifically, for each distant inertial observer at $\mathcal{I}^{+}$, a supertranslation corresponds to a simple change in the origin of the time coordinate. To understand why they are symmetries of future null infinity, we consider the following thought experiment.
Consider some asymptotically flat spacetime with an isolated astrophysical event—like a supernova explosion or a binary black hole merger—that is emitting radiation, such as photons or gravitational waves, outward in a spherical manner. Furthermore, imagine a network of distant inertial observers that are surrounding this event, each at some finite radius from the event that need not be the same as the other observers. If these observers can communicate, then they could—in principle—use their knowledge of their locations relative to each other and to the central event to synchronize their clocks to ensure that their measurements of the central event are simultaneous in some sense. Now, consider what happens if these observers were placed at larger (but finite) radii. At these new, farther away positions, the signals that they use to synchronize their clocks will take longer to travel between them, but there is no fundamental obstacle to this synchronization in principle.
@@ II.5 The effects of BMS transformations
Figure 1: “Cylinder” diagram of $\mathcal{I}^{+}$ to provide intuition about (i) boosts, (ii) space translations, and (iii) proper supertranslations. The retarded time coordinate $u$ runs vertically in each plot, while the polar coordinate $\theta$ runs azimuthally. Black circles correspond to surfaces of constant $u$ in the untransformed frame. A Lorentz boost dilates $u$ by some factor at each $\theta$ point, while a space translation and a proper supertranslation instead shift $u$ by some function of $\theta$, which can be written as some combination of $\ell=1$ spherical harmonics for space translations, and some combination of $\ell\geq 2$ spherical harmonics for proper supertranslations. The boost is in the $\hat{z}$ direction; the space translation is proportional to the spherical harmonic $Y_{(1,0)}(\theta,\phi)$, i.e., the $\hat{z}$ direction; and the proper supertranslation is proportional the spherical harmonic $Y_{(2,0)}(\theta,\phi)$.
@@
From this coordinate transformation, by examining how the coordinate tetrad on $\mathcal{I}^{+}$ transforms, one can then ascertain how asymptotic data on $\mathcal{I}^{+}$ that describes the spacetime metric transforms. But what is this data? Obviously, we want something on $\mathcal{I}^{+}$ that represents the gravitational wave strain $h$, since this is what our detectors measure. But is there other data that we should also be interested in, e.g., something describing the mass or the angular momentum of the spacetime that may be useful for, say, measuring the mass or the spin of an isolated black hole? While often neglected in the majority of the NR literature, the answer to this, as suggested by the isolated black hole example, is yes.
Apart from the strain, to fully reconstruct the metric on $\mathcal{I}^{+}$ one also needs other information about the spacetime, which is neatly encoded in the five complex Weyl scalars $\Psi_{0,1,2,3,4}$, i.e., the components of the Weyl tensor Newman and Penrose (1962). [^9] Note that even though these are called “scalars”, the Weyl scalars are also not invariant in any sense since they are still functions of the spacetime coordinates. The Weyl tensor measures the curvature of spacetime. However, the individual Weyl scalars each have their own unique interpretation. Specifically, they can be viewed as
  • $\Psi_{0}$: ingoing radiation;
  • $\Psi_{1}$: current multipole moment;
  • $\Psi_{2}$: mass multipole moment;
  • $\Psi_{3}$: news ($\sim\dot{h}$);
  • $\Psi_{4}$: outgoing radiation ($\sim\ddot{h}$),
where dots represent time derivatives. [^10] Note that other interpretations also exist, such as Ref. Szekeres (1965). Consequently, it can be seen that $\Psi_{3}$ and $\Psi_{4}$ are actually degenerate with with the strain and thus only the strain and $\Psi_{0,1,2}$ are needed to measure physical features of the spacetime. [^11] Formally, because of the Bianchi identities (see Eqs. (20)), one only needs the strain on $\mathcal{I}^{+}$ and $\Psi_{0,1,2}$ on a certain time slice of $\mathcal{I}^{+}$ to fully reconstruct the asymptotic spacetime metric As we will see later on in Sec. IV, $\Psi_{2}$ can be used to measure the mass of the spacetime and $\Psi_{1}$ can be used to measure the angular momentum of the spacetime.
If we now study how this asymptotic data transforms under a BMS transformation, one can ascertain through the tetrad transformation that the gravitational wave shear $\sigma$, which is related to the strain $h$ via $h\equiv 2\bar{\sigma}$, [^12] We introduce the shear because it makes many of the subsequent equations in this paper simpler; it is formally constructed in Eq. (V) by contracting complex dyads on the two-sphere with the angular part of the Bondi-Sachs metric. and the Weyl scalars transform as Bondi et al. (1962); Sachs (1962a); Newman and Unti (1962)
@@
When acting on spin-weighted spherical harmonics $\phantom{}{}_{s}Y_{(\ell,m)}$, they yield
ð$\eth\left(\phantom{}{}_{s}Y_{(\ell,m)}\right)=+\frac{1}{\sqrt{2}}\sqrt{(\ell-s)(\ell+s+1)}\phantom{}_{s+1}Y_{(\ell,m)},$ (12a)
ð¯$\bar{\eth}\left(\phantom{}{}_{s}Y_{(\ell,m)}\right)=-\frac{1}{\sqrt{2}}\sqrt{(\ell+s)(\ell-s+1)}\phantom{}_{s-1}Y_{(\ell,m)}.$ (12b)
@@
With this information, one then has everything that is needed to transform asymptotic data on $\mathcal{I}^{+}$ and therefore fix the coordinate freedom of that data to match some canonical frame. This notion of mapping asymptotic data to a certain frame is called BMS frame fixing and will be reviewed in the context of NR simulations in Sec. VII.
@@ II.6 Pedagogical approach to memory
This expression, in fact, is nearly correct. The one piece of information that is missing is that these two terms on the right-hand side of Eq. (II.6) need not fully cancel out. One way to realize this is by considering the scattering of two particles in linearized gravity Kovacs and Thorne (1978). First note that, because of the linearization, there will be no energy flux. However, because the particles scatter, there will still be a change in the angle-dependent mass, i.e., a change in the mass multipole moment. Then, because a change in the mass multipole moment corresponds to a change in the strain—or, equivalently, the shear—one can intuit that there should also be a term on the left-hand side of Eq. (II.6) that corresponds to the shear, e.g.,
@@
Using Eq. (II.6), one can then learn a large amount of interesting physics about gravitational waves. For one, we observe that, apart from the gravitational wave strain being sourced by the mass multipole moment, there is also a contribution from an angle-dependent energy flux. Furthermore, if we think about the net change in the strain between non-radiative regimes, i.e., the memory, we immediately find that there are two contributions: one from the angle-dependent mass and one from the angle-dependent energy. These unique contributions are the ordinary and null contributions to the memory Bieri and Garfinkle (2014). [^14] Previously, these contributions were called linear and nonlinear (or Christodoulou). However, they have since been renamed to better reflect the way in which they are sourced Zel’dovich and Polnarev (1974); Braginsky and Grishchuk (1985); Braginsky and Thorne (1987); Christodoulou (1991); Thorne (1992); Bieri and Garfinkle (2014). They can be understood as being sourced by two types of physical processes. In particular, ordinary memory is sourced by systems that contain unbound masses, i.e., massive particles approaching timelike infinity Bieri and Garfinkle (2014); Mädler and Winicour (2016); Zhang et al. (2024). Examples of systems that source this type of effect are hyperbolic black holes, neutron star disruptions, or even supernovae events. Meanwhile, null memory is sourced by null radiation that escapes to future null infinity, e.g., gravitational or electromagnetic waves. For an example of what the null memory looks like in the GW produced by a binary black hole merger, see Fig. 2.